# Adversarial Robustness in Two-Stage Learning-to-Defer: Algorithms and Guarantees

### v2

@@ E.2 Regression Task
We train three experts offline with three layers MLPs (128, 64, 32), each specializing in a specific subset of the dataset based on a predefined localization criterion. The first expert M1 trains on Southern California (latitude lower than 36), the second expert M2 on Central California (latitude between 36 and 38.5), and the last in Northern California (otherwise) representing a smaller area. MLPs are trained using a ReLU, an Adam optimizer (Kingma & Ba, 2017), a learning rate of $0.001$, and $500$ epochs. Agent costs are defined as $c_{0}(f(x),t)=\text{RMSE}(g(x),t)$ for the model and $c_{j>0}(m_{j}^{f}(x),t)=\text{RMSE}(m_{j}(x),t)$, consistent with (Mao et al., 2024c). We report respective RMSE of experts in Table 6.
@@
We train a regressor using a two-layer MLP with hidden dimensions (64, 32) on the full training set. The model uses ReLU activations, the Adam optimizer (Kingma & Ba, 2017), a learning rate of $0.001$, and is trained for 500 epochs. We report performance of the checkpoints corresponding to the lower empirical risk on the validation set. The model performance is reported in Table 6.
@@
Table 6: Agent RMSE on the California Housing validation set ($20$% of the dataset).
@@
We train a rejector using a MLP (8,16) for 100 epochs, a learning rate of $0.01$, a cosine scheduler, Adam optimizer (Kingma & Ba, 2017), and a batch size of 8096. We report performance of the checkpoints corresponding to the lower empirical risk on the validation set.
@@
We train a rejector using a MLP (8,16) for 400 epochs, a learning rate of $0.01$, a cosine scheduler, Adam optimizer (Kingma & Ba, 2017) with L2 weight decay of $10^{-4}$ acting as regularizer, and a batch size of $8096$. The hyperparameters are set to $\rho=1$ and $\nu=0.05$. The supremum component from the adversarial inputs is estimated using PGD10 (Madry et al., 2017) with $\epsilon$ equal to 25% of the variance of dataset’s features, the $\ell_{\infty}$ norm, and a step size of $\epsilon/10$, following the approach in (Mao et al., 2023b; Awasthi et al., 2023).
@@ E.3 Multi Task
We train two specialized experts using a Faster R-CNN (Ren et al., 2015) architecture with a MobileNet (Howard et al., 2017) backbone. The first expert, M1, is trained on images containing animals, while the second expert, M2, is trained on images containing vehicles. Both experts are trained using the Adam optimizer (Kingma & Ba, 2017) with a learning rate of $0.005$, a batch size of $128$, and trained for $50$ epochs. Agent costs are defined as $c_{0}(g(x),z)=\text{mAP}(g(x),z)$ for the model and $c_{j>0}(m_{j}(x),z)=\text{mAP}(m_{j}(x),z)$, consistent with (Montreuil et al., 2024). We report respective mAP of experts in Table 8.
@@
We train an object detection model using a larger Faster R-CNN (Ren et al., 2015) with ResNet-50 FPN (He et al., 2015) backbone. We train this model with Adam optimizer (Kingma & Ba, 2017), a learning rate of $0.005$, a batch size of $128$, and trained for $50$ epochs. We report performance of the checkpoints corresponding to the lower empirical risk on the validation set. The model performance is reported in Table 8.
@@
Baseline (Montreuil et al., 2024):
We train a rejector using a Faster R-CNN (Ren et al., 2015) with a MobileNet backbone (Howard et al., 2017) and a classification head. We train this rejector for $70$ epochs, a learning rate $5e^{-4}$, a cosine scheduler, Adam optimizer (Kingma & Ba, 2017), and a batch size of $256$. We report performance of the checkpoints corresponding to the lower empirical risk on the validation set.
@@
We train a rejector using a Faster R-CNN (Ren et al., 2015) with a MobileNet backbone (Howard et al., 2017) and a classification head for 70 epochs with a learning rate of $0.001$, a cosine scheduler, the Adam optimizer (Kingma & Ba, 2017) with L2 weight decay of $10^{-4}$ acting as regularizer, and a batch size of $64$. The hyperparameters are set to $\rho=1$ and $\nu=0.01$. The supremum component from the adversarial inputs is estimated using PGD20 (Madry et al., 2017) with $\epsilon=8/255$, the $\ell_{\infty}$ norm, and a step size of $\epsilon/20$, following the approach in (Mao et al., 2023b; Awasthi et al., 2023).
@@
Table 9: Performance comparison of SARD with the baseline (Montreuil et al., 2024) on the Pascal VOC dataset. The table reports mean Average Precision (mAP) under clean and adversarial scenarios.

### v3

@@ E.2 Regression Task
We train three experts offline with three layers MLPs (128, 64, 32), each specializing in a specific subset of the dataset based on a predefined localization criterion. The first expert M1 trains on Southern California (latitude lower than 36), the second expert M2 on Central California (latitude between 36 and 38.5), and the last in Northern California (otherwise) representing a smaller area. MLPs are trained using a ReLU, an Adam optimizer (Kingma & Ba, 2015), a learning rate of $0.001$, and $500$ epochs. Agent costs are defined as $c_{0}(f(x),t)=\text{RMSE}(g(x),t)$ for the model and $c_{j>0}(m_{j}^{f}(x),t)=\text{RMSE}(m_{j}(x),t)$, consistent with (Mao et al., 2024c). We report respective RMSE of experts in Table 6.
@@
We train a regressor using a two-layer MLP with hidden dimensions (64, 32) on the full training set. The model uses ReLU activations, the Adam optimizer (Kingma & Ba, 2015), a learning rate of $0.001$, and is trained for 500 epochs. We report performance of the checkpoints corresponding to the lower empirical risk on the validation set. The model performance is reported in Table 6.
@@
Table 6: Agent RMSE on the California Housing validation set (2020% of the dataset).
@@
We train a rejector using a MLP (8,16) for 100 epochs, a learning rate of $0.01$, a cosine scheduler, Adam optimizer (Kingma & Ba, 2015), and a batch size of 8096. We report performance of the checkpoints corresponding to the lower empirical risk on the validation set.
@@
We train a rejector using a MLP (8,16) for 400 epochs, a learning rate of $0.01$, a cosine scheduler, Adam optimizer (Kingma & Ba, 2015) with L2 weight decay of $10^{-4}$ acting as regularizer, and a batch size of $8096$. The hyperparameters are set to $\rho=1$ and $\nu=0.05$. The supremum component from the adversarial inputs is estimated using PGD10 (Madry et al., 2018) with $\epsilon$ equal to 25% of the variance of dataset’s features, the $\ell_{\infty}$ norm, and a step size of $\epsilon/10$, following the approach in (Mao et al., 2023b; Awasthi et al., 2023).
@@ E.3 Multi Task
We train two specialized experts using a Faster R-CNN (Ren et al., 2015) architecture with a MobileNet (Howard et al., 2017) backbone. The first expert, M1, is trained on images containing animals, while the second expert, M2, is trained on images containing vehicles. Both experts are trained using the Adam optimizer (Kingma & Ba, 2015) with a learning rate of $0.005$, a batch size of $128$, and trained for $50$ epochs. Agent costs are defined as $c_{0}(g(x),z)=\text{mAP}(g(x),z)$ for the model and $c_{j>0}(m_{j}(x),z)=\text{mAP}(m_{j}(x),z)$, consistent with (Montreuil et al., 2025c). We report respective mAP of experts in Table 8.
@@
We train an object detection model using a larger Faster R-CNN (Ren et al., 2015) with ResNet-50 FPN (He et al., 2015) backbone. We train this model with Adam optimizer (Kingma & Ba, 2015), a learning rate of $0.005$, a batch size of $128$, and trained for $50$ epochs. We report performance of the checkpoints corresponding to the lower empirical risk on the validation set. The model performance is reported in Table 8.
@@
Baseline (Montreuil et al., 2025c):
We train a rejector using a Faster R-CNN (Ren et al., 2015) with a MobileNet backbone (Howard et al., 2017) and a classification head. We train this rejector for $70$ epochs, a learning rate $5e^{-4}$, a cosine scheduler, Adam optimizer (Kingma & Ba, 2015), and a batch size of $256$. We report performance of the checkpoints corresponding to the lower empirical risk on the validation set.
@@
We train a rejector using a Faster R-CNN (Ren et al., 2015) with a MobileNet backbone (Howard et al., 2017) and a classification head for 70 epochs with a learning rate of $0.001$, a cosine scheduler, the Adam optimizer (Kingma & Ba, 2015) with L2 weight decay of $10^{-4}$ acting as regularizer, and a batch size of $64$. The hyperparameters are set to $\rho=1$ and $\nu=0.01$. The supremum component from the adversarial inputs is estimated using PGD20 (Madry et al., 2018) with $\epsilon=8/255$, the $\ell_{\infty}$ norm, and a step size of $\epsilon/20$, following the approach in (Mao et al., 2023b; Awasthi et al., 2023).
@@
Table 9: Performance comparison of SARD with the baseline (Montreuil et al., 2025c) on the Pascal VOC dataset. The table reports mean Average Precision (mAP) under clean and adversarial scenarios.